OpenStudy (anonymous):

NEED HELP NOW WILL FAN Line M is represented by the following equation: x + y = −1 What is most likely the equation for line P so the set of equations has infinitely many solutions?

OpenStudy (anonymous):

is there a grafe btw im not the best speller

OpenStudy (anonymous):

nope no graph

OpenStudy (anonymous):

the whole question is this Line M is represented by the following equation: x + y = −1 What is most likely the equation for line P so the set of equations has infinitely many solutions?

OpenStudy (anonymous):

so how would you know what line p is

OpenStudy (anonymous):

answers are What is most likely the equation for line P so the set of equations has infinitely many solutions? 2x + 2y = 2 2x + 2y = 4 2x + 2y = −2 x − y = 1

OpenStudy (anonymous):

whats line p

OpenStudy (anonymous):

thats what were looking for...

OpenStudy (anonymous):

is there some pic for this

OpenStudy (anonymous):

no i will post the whole question look Question 5 (Multiple Choice Worth 4 points) (08.01) Line M is represented by the following equation: x + y = −1 What is most likely the equation for line P so the set of equations has infinitely many solutions? 2x + 2y = 2 2x + 2y = 4 2x + 2y = −2 x − y = 1